# takeawaybox backgroundcolor=gray!30, shadow=false, innerleftmargin=5pt, innerrightmargin=5pt, roundcorner=10pt, skipabove=12pt

### v3

@@ 1 Introduction
To answer these questions we conduct a thorough literature review, collecting detailed information from $>$130 papers. Although previous surveys on Mechanistic Interpretability (Ferrando et al., 2024; Bereska & Gavves, 2024) and frameworks for causal interpretability (Mueller et al., 2024; Geiger et al., 2024a) have touched on RepE, there is no dedicated survey of this research area. A survey is especially pressing since over 100 papers have been released within the last year without a systematization of the literature. In addition to Activation Steering, we include methods that modify the weights, identify concept representations through optimizing for outputs, and are not based on vectors. We focus on methods that apply control on intermediate representations of a model without fully replacing them, thus excluding soft-prompting, decoding-based methods, and Activation Patching. For a detailed description of the inclusion criteria and literature search process, see Appendix B.
@@ 2 What is Representation Engineering?
To achieve this, a pipeline with two steps is necessary: First, we identify how the targeted concept is represented within the model. Secondly, we use that information to steer the model’s representations on new inputs. In this context, a concept can be understood broadly as any human-understandable feature like behaviors, tasks, entities, pieces of knowledge, or personality traits. RepE assumes that such concepts are represented in LLMs. Representations are the internal structures or encodings that a model uses to capture and process information about a concept.
@@
2.1 Illustrative Examples
To illustrate RepE, we provide an example inspired by Zou et al. (2023a), where RepE is applied to make an LLM give more truthful answers. To identify how truthfulness is represented, inputs are provided that instruct the model to be honest or to lie. Next, the activations for these inputs are collected. Then, the difference between activations for honest and dishonest inputs is calculated. This produces a vector that denotes a direction in the activation space which represents honesty. This vector can be used to guide the training of a LoRA module that generates activations more aligned with the truthfulness direction. This effectively steers the model to give more truthful answers.
@@ 2.2 Advantages of Representation Engineering
{mdframed}
[style=takeawaybox] Takeaway: What is Representation Engineering?
RepE manipulates the representations of a model in order to control its behavior w.r.t. a concept. This is done by identifying how a concept is represented within the model and then using that knowledge to steer the representations. It allows for effective, efficient, flexible, and precise control over models’ behavior while largely retaining quality.
@@ 3.1 Representation Identification
Methods for Representation Identification need to carry out three steps. First, they need to provide a stimulus related to the concept. These are strings used as inputs or ground-truth outputs that have a positive, negative, neutral, or unknown relationship to the concept. This means that they exemplify or contain the concept, serve as counterexamples to the concept, are unrelated to the concept, or that their relationship to the concept is unknown. This relationship is usually predefined by humans, e.g., as labels in a dataset. However, in some cases, it is evaluated dynamically during the execution of the method. It is important that these strings actually activate the desired concept’s representation, so it is present in the activations. Secondly, they collect activations of the model related to these strings.
@@
The activation collection function $Act$
@@ 3.3 Representation Control
{mdframed}
[style=takeawaybox] Takeaway: Framework & Notation
@@
  1. 1.
    Representation Identification finds a concept operator that denotes the model’s representation of the concept.
  2. 2.
    Representation Control applies a steering function that uses the concept operator to modify the activations or weights of the model.
Further, RepE pipelines make assumptions about representations to operationalize them into a concept operator.
@@ 4.1 Input Reading
RI methods that fall into the Input Reading category provide inputs related to the concept, read out their activations, and finally calculate the operator from the resulting sets of activations.
@@ 4.1.3 Calculating the Concept Operator
Difference in Means. Taking the difference between the classes of activations can identify the direction that captures their differences. Difference-in-Means (DiM) is a popular method that simply calculates the average for both classes and subtracts the negative from the positive one (Panickssery et al., 2024; Jorgensen et al., 2024; Ball et al., 2024). This is equivalent to taking the average of the difference between pairs of examples $o^{c}=\frac{1}{n}\sum_{a^{+},a^{-}\in A^{+},A^{-}}a^{+}-a^{-}$ Earlier papers took the difference between a single pair of inputs (Turner et al., 2024; Liu et al., 2023), however, it has been found that taking the difference across larger datasets leads to higher-quality operators (Chu et al., 2024; Jorgensen et al., 2024). Arora et al. (2024) employ an unsupervised method that discovers 2 clusters of activations through $k$-means and subtracts their means. Im & Li (2025) show that DiM is optimal at identifying a vector that matches negative examples to positive ones. However, these vectors are not suited to steering positive examples where they lead to large performance reductions.
@@ 4.2 Output Optimization
Representation Identification methods that perform Output Optimization provide information about the concept by scoring outputs. This is done by first providing inputs and collecting their activations and outputs. Secondly, a loss function is constructed by evaluating the outputs with regard to the concept. Thirdly, an optimization algorithm is used to iteratively optimize the activations to achieve a high concept score.
@@ 4.3 Unsupervised Feature Learning
Representation Identification methods that perform Unsupervised Feature Learning (UFL) first learn sparse features in the model’s activations and then identify which concepts they belong to. This is done by first collecting a set of activations, applying a learning algorithm that discovers features in the activation space, and then identifying which human-understandable concept a feature corresponds to. In contrast to other RI approaches, UFL methods do not set out to find one specific concept representation but identify a large set of concepts from which concepts of interest can be chosen.
@@ 4.3.4 Applying SAEs
{mdframed}
[style=takeawaybox] Takeaway: Representation Identification
@@
  1. 1.
    Input Reading: Construct sets of inputs related to the concept, collect their activations, and calculate the concept operator from the sets of activations.
  2. 2.
    Output Optimization: Collects activations and outputs, scores the outputs according to the concept, and optimizes a concept operator for that score.
  3. 3.
    Unsupervised Feature Learning: Collect activations, discover features in the activations with an unsupervised learning algorithm, and identify which concepts the features correspond to.
@@ 5.1 Assumed Geometry
There is an ongoing discussion in the interpretability community about the geometry that transformer models use to represent internal features. The Linear Representation Hypothesis poses that concepts are represented as linear directions in the activation space, and their intensity is denoted by the magnitude of the activations in that direction (Park et al., 2024b). Although this hypothesis has some merits and, if true, would make RepE easier, it has also been criticized (see section 11 for discussion). Thus, RepE methods with alternative feature geometries have been proposed. They assume representations to be fuzzy, clusters (Postmus & Abreu, 2024), or non-linear manifolds (Qiu et al., 2024). Other possibilities for feature geometries include circles (Engels et al., 2025), onions (Csordás et al., 2024), non-linear paths, or subsets of a vector space.
@@ 5.2 Concept Operators
{mdframed}
[style=takeawaybox] Takeaways: Representation Operationalization
  1. 1.
    Representaiton Geometry: RepE methods make implicit assumptions about the geometry of representations. Many assume that concepts are represented as linear directions, but alternative perspectives exist.
  2. 2.
    Concept Operator: Most commonly, the concept operator is a vector denoting a linear direction. Others use matrices to capture principled directions, covariances or all linearly available information about a concept.
@@ 6.1 Modifying Activations
Activations can be controlled at inference time with an activation steering function $F_{a}(a_{l},o^{c}_{l})\rightarrow a_{l}^{c}$ (see Definition 4). For a given input, we read the activations at an intermediate layer $a_{l}$ during the forward pass, perform some operation $F$ on it using the operator $o^{c}_{l}$, and then use the modified activations $a^{c}_{l}$ as input for the next layer. By steering the activations, we can control the model’s output with regard to the concept.
@@ 6.1.1 Activation Steering Functions
Multiple papers attempt to add multiple steering vectors at once to steer multiple concepts or multiple aspects of one concept. This can be done by simply adding all vectors at once (Wang et al., 2024c; Scalena et al., 2024) or combining them in a weighted sum (Chen et al., 2024b; Zeng et al., 2024). However, this leads to interference between concepts and increases negative impact on the LLM’s quality (see subsection 12.1). van der Weij et al. (2024) avoid this by adding different vectors at different layers. Adila et al. (2024b) decompose the different vectors into their orthonormal basis and use those for steering.
@@ 6.2 Modifying Weights
The representations of a model can be steered by modifying the weights with a weight steering function $F^{w}(w_{l},o^{c}_{l})\rightarrow w_{l}^{c}$.
@@ 6.2.2 Editing Weights
{mdframed}
[style=takeawaybox] Takeaway: Representation Control
@@
  1. 1.
    Activations can be steered during inference by addition, rejection, or projection, thus increasing, removing, or adapting the concept representation. Activation steering can be applied at different tokens and dynamically adjusted for new inputs.
  2. 2.
    Weights can be steered by training them to generate desired activations and outputs or by editing components of weight matrices.
@@ 7.1 Prototypical RepE Pipelines
In Table 7.1, we classify RepE pipelines according to which strategy they employ for Representation Identification, the shape of their operator, and whether they control activations or weights. We score Representation Identification methods on whether they require a dataset labeled according to the concept and whether they require expensive optimization. Additionally, we score whether the method is restricted to linear representations or can capture non-linear ones. Furthermore, we score whether controlling the model requires retraining of its weights and whether the operation to control the model is dependent on the given inputs. Lastly, we determine the types of concepts the method was initially designed to engineer. However, this does not mean that they are not able to engineer other concepts. We select these RepE pipelines to show a diversity of approaches while preferring highly-cited papers that were published at top conferences.
@@
Method Representation Identification Representations Operationalization Representation Control Originally proposed for
Name Type Works with unlabeled dataset? Optimi- zation Free? Concept Operator Allows non-linear representations? Type Training- free? Context- adjusted steering?
[…]
@@ 7.2.1 What Works for Representation Identification?
Is it more effective to use Input Reading or Output Optimization? - Depends on the concept. Multiple papers claim that their method based on output optimization improves on input reading methods (Yin et al., 2024; Cai et al., 2024; Cao et al., 2024a; Ackerman, 2024). However, we lack evidence for the contrary, which is likely because papers that propose input reading techniques did not compare their methods to output optimization methods. Two unbiased comparisons find that an input reading method is better at steering the sentiment of text (Konen et al., 2024) and that output optimization methods control the concrete topics more effectively (Wu et al., 2025).
@@
Are steering features from Sparse Auto-Encoders better than other RI methods? - Possibly a combination. In Zhao et al. (2024) and Makelo et al. (2025), an SAE-based steering is more effective than input-reading-based RI methods. In contrast,Chalnev et al. (2024) find that an input-reading-based RI method leads to more effective steering than SAE features. However, results in Chalnev et al. (2024) and Kharlapenko et al. (2024) suggest that SAEs can be used to make input-reading-based concept operators more effective. Finally, evaluations on the AxBench (Wu et al., 2025) find that SAEs perform worse at concept detection and steering than LoReFT and class means but better than methods using probing or dimensionality reduction.
@@
Which prompting format works best? - Unclear. Braun et al. (2025) compare control effectiveness for seven prompt settings varying in whether an answer token is appended, whether an instruction is prepended, and whether few-shot demonstrations are included. While the setting that only appends an answer token has a slightly larger average control effectiveness, they conclude that no prompt type clearly outperforms the other. Wang & Shu (2023) find that choice prompts, where the model has to select answer A or B, outperform free-form answers. This is likely because the choice prompt concentrates the relevant context into a single token and thus more effectively triggers the concept representation. The lack of principled evaluations of optimal prompting formats for eliciting representations makes this an important area of further study.
@@
At which token positions should activations be read out? - Unclear. Some evidence suggests that averaging the activations of multiple tokens is better than just using the last token’s activations (Hoscilowicz et al., 2024). Furthermore, focusing on the last token given by the user can be more effective than taking activations from an appended post-prompt like “I think the {attribute} of this user is” (Chen et al., 2024a). Braun et al. (2025) do not find a clear difference between reading activations at the last token of the question or at the token containing the answer.
@@
Which function is best for calculating the difference between sets of activations? - Difference-in-Means. In Input Reading, functions like Difference-in-Means (DiM), linear probes, Principle Component Analysis, or Contrast Consistent Search (Burns et al., 2023) are used to identify the concept representation from contrasting sets of activations. This has been experimentally studied in 7 papers (Li et al., 2023a; Marks & Tegmark, 2023; von Rütte et al., 2024; Xu et al., 2024a; Arora et al., 2024; Wu et al., 2025; Im & Li, 2025). In 6 out of 7 papers Difference-in-Means leads to the strongest performance. The next best seems to be taking the weight vector of a linear probe. PCA and CCS tend to perform worse. A fair comparison between DiM, PCA, and classifier-based methods finds that DiM is best at positively and negatively steering multiple-choice questions, equals the performance of probes at steering long-form generations, but leads to the largest performance deterioration on positive examples (Im & Li, 2025). Lastly, Distributed Alignment Search (DAS) (Geiger et al., 2024b) was only tested once in which it performed the strongest, highlighting that DAS remains underexplored. Besides Input Reading, it is advantageous to apply mean-centering to the steered activations as shown by Jorgensen et al. (2024).

### v4

# Taxonomy, Opportunities, and Challenges of Representation Engineering for Large Language Models

@@ 1 Introduction
  • How are RepE methods evaluated? We describe evaluation methodologies and benchmarks, and propose best practices for evaluating RepE methods.
@@
To answer these questions we conduct a thorough literature review, collecting detailed information from >>130 papers. Although previous surveys on Mechanistic Interpretability (Ferrando et al., 2024; Bereska & Gavves, 2024) and frameworks for causal interpretability (Mueller et al., 2024; Geiger et al., 2024a) have touched on RepE, there is no dedicated survey of this research area. A survey is especially pressing since over 100 papers have been released within the last year without a systematization of the literature. In addition to Activation Steering, we include methods that modify the weights, identify concept representations through optimizing for outputs, and are not based on vectors. We focus on methods that apply control on intermediate representations of a model without fully replacing them, thus excluding soft-prompting, decoding-based methods, and Activation Patching. For a detailed description of the inclusion criteria and literature search process, see Appendix B.
@@ 2 What is Representation Engineering?
One perspective on RepE frames the computations in an LLMs as a computer program with intermediate variables and computations that use these variables. In this perspective we can attempt to find intermediate variable that correspond to specific concepts. This allows us to meaningfully change the values of these variables to influence later computations and ultimately the outputs of the program. Another perspective states that RepE aims to identify patterns in the activations of an LLM that correspond to high-level, human-understandable concepts. It turns out that these activation patterns can be used to manipulate the activations to get consistent changes in the behavior of the LLM.
Goals of RepE. The goals of RepE are (1) Behavior Steering: Our ability to control the behavior of LLMs and (2) Interpretability: Our ability to understand the internal computations of LLMs. RepE can be used by LLM providers to prevent undesired behavior of an LLM such as guarding against jailbreaks. It can also be used to finely control the behavioral tendencies of the LLM or to improve it’s suitability and performance at specific tasks for example by adapting it’s personality and style or by improving it’s reasoning performance. Additionally, RepE can provide us with some insight into the internal processes of an LLM which are commonly considered as a black-box. By demonstrating that some concepts are consistently represented in LLMs, determining in how much a concept’s representation is active on specific outputs or showcasing the influence that a concept has on the output of the model. As such RepE is often a tool for scientific inquiry into specific LLM behaviors.
Running Example. One of the most commonly steered concepts is Truthfulness, a models intention to say things that are true instead of false as for example measured in the TruthfulQA benchmark. RepE methods aim to identify patterns in the activations that correspond to the models representation of the concept. This could for example be a linear direction whose magnitude denotes how truthful the outputs of an LLM are on a specific inputs. This activation pattern, hereforth called concept operator, can then be used to detect when the model is being dishonest and to steer it to give truthful outputs. As such it could be deployed by model providers to decrease hallucinations and be used by safety researchers to detect and mitigate cases of deceptive behavior.
The RepE Pipeline. RepE methods consist of a pipeline with two steps: First, we identify how the targeted concept is represented within the model. Secondly, we use that information to steer the model’s representations on new inputs. In this context, a concept can be understood broadly as any human-understandable feature like behaviors, tasks, entities, pieces of knowledge, or personality traits. RepE assumes that such concepts are represented in LLMs. Representations are the internal structures or encodings that a model uses to capture and process information about a concept.
@@
2.1 Strategies for RepE
There are a variety of techniques to perform Representation Identification and Control. The most common way to perform RepE is through Linear Activation Addition. Hereby, a linear direction is identified that should denote the representation of a concept for example by contrasting activations for inputs where the concept is or isn’t present. Another popular approach are Sparse Autoencoders, that disentangle features into a dictionary of concepts by projecting them into a higher-dimensional, sparse space. Finally, it’s possible to directly identify representations that cause the specific behavior changes by fine-tuning a concept operator towards desired outputs.
Once the representation was identified, we can use it to control the LLMs behavior by adapting it’s activations or changing it’s weights. For example we could simply add a vector on the activations for new inputs. Or we can adapt the weights, such that they produce activations that align more strongly with the activation patterns we have identified.
Illustrative Examples To illustrate RepE, we provide an example inspired by Zou et al. (2023a), where RepE is applied to make an LLM give more truthful answers. To identify how truthfulness is represented, inputs are provided that instruct the model to be honest or to lie. Next, the activations for these inputs are collected. Then, the difference between activations for honest and dishonest inputs is calculated. This produces a vector that denotes a direction in the activation space which represents honesty. This vector can be used to guide the training of a LoRA module that generates activations more aligned with the truthfulness direction. This effectively steers the model to give more truthful answers.
@@ 2.2 Advantages of Representation Engineering
Takeaway: What is Representation Engineering?
RepE manipulates the representations of a model in order to control its behavior w.r.t. a concept. It can be used to steer a models behavior and interpret its internals. This is done by identifying how a concept is represented within the model and then using that knowledge to steer the representations. It allows for effective, efficient, flexible, and precise control over models’ behavior while largely retaining quality.
@@ 3.1 Representation Identification
Methods for Representation Identification need to carry out three steps. First, they need to provide a stimulus related to the concept. These are strings used as inputs or ground-truth outputs that have a positive, negative, neutral, or unknown relationship to the concept. This means that they exemplify or contain the concept, serve as counterexamples to the concept, are unrelated to the concept, or that their relationship to the concept is unknown. In our running example these strings might be truthful or deceptive outputs, text that is neither honest nor dishonest or we might just not know it’s relationship to honesty. This relationship is usually predefined by humans, e.g., as labels in a dataset. However, in some cases, it is evaluated dynamically during the execution of the method. It is important that these strings actually activate the desired concept’s representation, so it is present in the activations. Secondly, they collect activations of the model related to these strings.
@@
The activation collection function ActAct
@@ 3.3 Representation Control
Takeaway: Framework & Notation
@@
1. Representation Identification finds a concept operator that denotes the model’s representation of the concept. 2. Representation Control applies a steering function that uses the concept operator to modify the activations or weights of the model. Further, RepE pipelines make assumptions about representations to operationalize them into a concept operator.
@@ 4.1 Input Reading
RI methods that fall into the Input Reading category provide inputs related to the concept, read out their activations, and finally calculate the operator from the resulting sets of activations. In our running example we can feed inputs related to honest or dishonest behavior into the model, collect their activations and identify patterns that correspond to honesty.
@@ 4.1.3 Calculating the Concept Operator
Difference in Means. Taking the difference between the classes of activations can identify the direction that captures their differences. Difference-in-Means (DiM) is a popular method that simply calculates the average for both classes and subtracts the negative from the positive one (Panickssery et al., 2024; Jorgensen et al., 2024; Ball et al., 2024). This is equivalent to taking the average of the difference between pairs of examples $o^{c}=\frac{1}{n}\sum_{a^{+},a^{-}\in A^{+},A^{-}}a^{+}-a^{-}$ . Earlier papers took the difference between a single pair of inputs (Turner et al., 2024; Liu et al., 2023), however, it has been found that taking the difference across larger datasets leads to higher-quality operators (Chu et al., 2024; Jorgensen et al., 2024). Arora et al. (2024) employ an unsupervised method that discovers 2 clusters of activations through $k$-means and subtracts their means. Im & Li (2025) show that DiM is optimal at identifying a vector that matches negative examples to positive ones. However, these vectors are not suited to steering positive examples where they lead to large performance reductions.
@@ 4.2 Output Optimization
Representation Identification methods that perform Output Optimization provide information about the concept by scoring outputs. This is done by first providing inputs and collecting their activations and outputs. Secondly, a loss function is constructed by evaluating the outputs with regard to the concept. Thirdly, an optimization algorithm is used to iteratively optimize the activations to achieve a high concept score. In our running example this can be done by defining a way to score whether model generations are truthful and then optimizing a concept operator that can be applied to make truthful outputs more likely.
@@ 4.3 Unsupervised Feature Learning
Representation Identification methods that perform Unsupervised Feature Learning (UFL) first learn sparse features in the model’s activations and then identify which concepts they belong to. This is done by first collecting a set of activations, applying a learning algorithm that discovers features in the activation space, and then identifying which human-understandable concept a feature corresponds to. In contrast to other RI approaches, UFL methods do not set out to find one specific concept representation but identify a large set of concepts from which concepts of interest can be chosen. In our running example, we would first discover a library of representations for many concepts and then identify which one corresponds to truthfulness.
@@ 4.3.4 Applying SAEs
Takeaway: Representation Identification
@@
1. Input Reading: Construct sets of inputs related to the concept, collect their activations, and calculate the concept operator from the sets of activations. 2. Output Optimization: Collects activations and outputs, scores the outputs according to the concept, and optimizes a concept operator for that score. 3. Unsupervised Feature Learning: Collect activations, discover features in the activations with an unsupervised learning algorithm, and identify which concepts the features correspond to.
@@ 5.1 Assumed Geometry
There is an ongoing discussion in the interpretability community about the geometry that transformer models use to represent internal features. The Linear Representation Hypothesis poses that concepts are represented as linear directions in the activation space, and their intensity is denoted by the magnitude of the activations in that direction (Park et al., 2024b). Although this hypothesis has some merits and, if true, would make RepE easier, it has also been criticized (see Section 12 for discussion). Thus, RepE methods with alternative feature geometries have been proposed. They assume representations to be fuzzy, clusters (Postmus & Abreu, 2024), or non-linear manifolds (Qiu et al., 2024). Other possibilities for feature geometries include circles (Engels et al., 2025), onions (Csordás et al., 2024), non-linear paths, or subsets of a vector space. See Section 13.2.2 for consequences of these assumptions and Section 14.2 for possibilities for extension.
@@ 5.2 Concept Operators
Takeaways: Representation Operationalization 1. Representaiton Geometry: RepE methods make implicit assumptions about the geometry of representations. Many assume that concepts are represented as linear directions, but alternative perspectives exist. 2. Concept Operator: Most commonly, the concept operator is a vector denoting a linear direction. Others use matrices to capture principled directions, covariances or all linearly available information about a concept.
@@ 6.1 Modifying Activations
Activations can be controlled at inference time with an activation steering function $F_{a}(a_{l},o^{c}_{l})\rightarrow a_{l}^{c}$ (see Definition 4). For a given input, we read the activations at an intermediate layer $a_{l}$ during the forward pass, perform some operation $F$ on it using the operator $o^{c}_{l}$, and then use the modified activations $a^{c}_{l}$ as input for the next layer. By steering the activations, we can control the model’s output with regard to the concept. In our running example, we would take activations and steer them to be more similar to truthful activations and less similar to dishonest ones, thus causing later computations and the outputs to be more honest.
@@ 6.1.1 Activation Steering Functions
Multiple papers attempt to add multiple steering vectors at once to steer multiple concepts or multiple aspects of one concept. This can be done by simply adding all vectors at once (Wang et al., 2024c; Scalena et al., 2024) or combining them in a weighted sum (Chen et al., 2024b; Zeng et al., 2024). However, this leads to interference between concepts and increases negative impact on the LLM’s quality (see Section 13.1). van der Weij et al. (2024) avoid this by adding different vectors at different layers. Adila et al. (2024b) decompose the different vectors into their orthonormal basis and use those for steering.
@@ 6.2 Modifying Weights
The representations of a model can be steered by modifying the weights with a weight steering function $F^{w}(w_{l},o^{c}_{l})\rightarrow w_{l}^{c}$. In our running example we would modify the weights of the model so that it produces activation patterns and outputs that are aligned with honesty.
@@ 6.2.2 Editing Weights
Takeaway: Representation Control
@@
1. Activations can be steered during inference by addition, rejection, or projection, thus increasing, removing, or adapting the concept representation. Activation steering can be applied at different tokens and dynamically adjusted for new inputs. 2. Weights can be steered by training them to generate desired activations and outputs or by editing components of weight matrices.
@@ 7.1 Prototypical RepE Pipelines
In Table 8, we classify RepE pipelines according to which strategy they employ for Representation Identification, the shape of their operator, and whether they control activations or weights. We score Representation Identification methods on whether they require a dataset labeled according to the concept and whether they require expensive optimization. Additionally, we score whether the method is restricted to linear representations or can capture non-linear ones. Furthermore, we score whether controlling the model requires retraining of its weights and whether the operation to control the model is dependent on the given inputs. Lastly, we determine the types of concepts the method was initially designed to engineer. However, this does not mean that they are not able to engineer other concepts. We select these RepE pipelines to show a diversity of approaches while preferring highly-cited papers that were published at top conferences.
@@
Method Representation Identification Representations Operationalization Representation Control
Originally
proposed for
Name Type
Works with
unlabeled
dataset?
Optimi-
zation
Free?
Concept
Operator
Allows
non-linear
representations?
Type
Training-
free?
Context-
adjusted
steering?
CAA (Panickssery et al., 2024) Input Reading q Vector q Activations q High-level behaviors
LoRRA (Zou et al., 2023a) Input Reading Vector q Weights q q High-level behaviors
ITI (Li et al., 2023a) Input Reading q q Vectors q Activations q Truthfulness
SAE (Templeton et al., 2024) Unsupervised Feature Learning q Vector q Activations q Abstract features
LoReFT (Wu et al., 2024c) Output Optimization q 2 Matrices + vector q Activations/ Weights q Improving task performance
BiPO (Cao et al., 2024a) Output Optimization q q Vector q Activations q Aligning to preferences
SEA (Qiu et al., 2024) Input Reading q 2 matrices Activations High-level behaviors
MiMiC Singh et al. (2024) Input Reading q Matrix + vector q Activations q Removing bias & Toxicity
@@ 7.2.1 What Works for Representation Identification?
Is it more effective to use Input Reading or Output Optimization? - Depends on the concept.
@@
Multiple papers claim that their method based on output optimization improves on input reading methods (Yin et al., 2024; Cai et al., 2024; Cao et al., 2024a; Ackerman, 2024). However, we lack evidence for the contrary, which is likely because papers that propose input reading techniques did not compare their methods to output optimization methods. Two unbiased comparisons find that an input reading method is better at steering the sentiment of text (Konen et al., 2024) and that output optimization methods control the concrete topics more effectively (Wu et al., 2025).
Are steering features from Sparse Auto-Encoders better than other RI methods? - Possibly a combination.
@@
In Zhao et al. (2024) and Makelo et al. (2025), an SAE-based steering is more effective than input-reading-based RI methods. In contrast,Chalnev et al. (2024) find that an input-reading-based RI method leads to more effective steering than SAE features. However, results in Chalnev et al. (2024) and Kharlapenko et al. (2024) suggest that SAEs can be used to make input-reading-based concept operators more effective. Finally, evaluations on the AxBench (Wu et al., 2025) find that SAEs perform worse at concept detection and steering than LoReFT and class means but better than methods using probing or dimensionality reduction.
Which prompting format works best? - Unclear.
Braun et al. (2025)
@@
compare control effectiveness for seven prompt settings varying in whether an answer token is appended, whether an instruction is prepended, and whether few-shot demonstrations are included. While the setting that only appends an answer token has a slightly larger average control effectiveness, they conclude that no prompt type clearly outperforms the other. Wang & Shu (2023) find that choice prompts, where the model has to select answer A or B, outperform free-form answers. This is likely because the choice prompt concentrates the relevant context into a single token and thus more effectively triggers the concept representation. The lack of principled evaluations of optimal prompting formats for eliciting representations makes this an important area of further study.
At which token positions should activations be read out? - Unclear.
@@
Some evidence suggests that averaging the activations of multiple tokens is better than just using the last token’s activations (Hoscilowicz et al., 2024). Furthermore, focusing on the last token given by the user can be more effective than taking activations from an appended post-prompt like “I think the {attribute} of this user is” (Chen et al., 2024a). Braun et al. (2025) do not find a clear difference between reading activations at the last token of the question or at the token containing the answer.
Which function is best for calculating the difference between sets of activations? - Difference-in-Means.
@@
In Input Reading, functions like Difference-in-Means (DiM), linear probes, Principle Component Analysis, or Contrast Consistent Search (Burns et al., 2023) are used to identify the concept representation from contrasting sets of activations. This has been experimentally studied in 7 papers (Li et al., 2023a; Marks & Tegmark, 2023; von Rütte et al., 2024; Xu et al., 2024a; Arora et al., 2024; Wu et al., 2025; Im & Li, 2025). In 6 out of 7 papers Difference-in-Means leads to the strongest performance. The next best seems to be taking the weight vector of a linear probe. PCA and CCS tend to perform worse. A fair comparison between DiM, PCA, and classifier-based methods finds that DiM is best at positively and negatively steering multiple-choice questions, equals the performance of probes at steering long-form generations, but leads to the largest performance deterioration on positive examples (Im & Li, 2025). Lastly, Distributed Alignment Search (DAS) (Geiger et al., 2024b) was only tested once in which it performed the strongest, highlighting that DAS remains underexplored. Besides Input Reading, it is advantageous to apply mean-centering to the steered activations as shown by Jorgensen et al. (2024).